# RooseBERT: A New Deal For Political Language Modelling

### v1

@@ Abstract
The increasing amount of political debates and politics-related discussions calls for the definition of novel computational methods to automatically analyse such content with the final goal of lightening up political deliberation to citizens. However, the specificity of the political language and the argumentative form of these debates (employing hidden communication strategies and leveraging implicit arguments) make this task very challenging, even for current general-purpose pre-trained Language Models. To address this issue, we introduce a novel pre-trained Language Model for political discourse language called RooseBERT. Pre-training a language model on a specialised domain presents different technical and linguistic challenges, requiring extensive computational resources and large-scale data. RooseBERT has been trained on large political debate and speech corpora (8K debates, each composed of several sub-debates on different topics) in English. To evaluate its performances, we fine-tuned it on four downstream tasks related to political debate analysis, i.e., named entity recognition, sentiment analysis, argument component detection and classification, and argument relation prediction and classification. Our results demonstrate significant improvements over general-purpose Language Models on these four tasks, highlighting how domain-specific pre-training enhances performance in political debate analysis. We release the RooseBERT language model for the research community.
Codehttps://github.com/deborahdore/RooseBERT
@@ 1 Introduction
Political debates, politics-related (online) discussions and opinions represent key elements of public discourse and deliberation, and they are fundamental for citizens to take informed decisions. The ability to automatically analyse such a kind of content is therefore central in Artificial Intelligence, where different aspects of political discourse have been investigated, e.g., performance evaluation (Liu et al. 2024), argumentative structures (Haddadan, Cabrio, and Villata 2019), sentiment analysis (Abercrombie and Batista-Navarro 2020), and logical fallacies (Goffredo et al. 2022). These approaches rely on Pre-trained Language Models (PLMs) originally developed for general-purpose corpora, e.g., BERT (Devlin et al. 2019). Although these models achieve good results in some of these tasks, they underperform when the goal is to capture the unique features of political content (e.g., domain-specific terminology, implicit argumentation, strategic communication patterns). In the literature, researchers have shown that domain-specific variations in vocabulary and semantics can significantly affect model performance (Nooralahzadeh, Øvrelid, and Lønning 2018), avoiding vocabulary mismatches and encoding semantic shifts (Lee et al. 2020).
To address this issue, in this paper, we introduce RooseBERT, a domain-specific language model pre-trained on English political debates. Built on the BERT architecture, widely adopted for domain adaptation tasks (Beltagy, Lo, and Cohan 2019; Huang, Altosaar, and Ranganath 2019; Lee et al. 2020; Yang, Uy, and Huang 2020; Kawintiranon and Singh 2022; Yang et al. 2023; Wada et al. 2024; Zeinali et al. 2024), RooseBERT is tailored to capture the distinctive features of political debates. Our aim is to improve the performance of state-of-the-art models on tasks such as sentiment analysis, stance detection, and argument mining which are especially relevant to political analysis. Encoder-based architectures like BERT are especially well-suited for such tasks due to their strength in encoding semantic relationships and contextual information and have proven strong performance on a range of downstream tasks relevant to political discourse analysis (Abercrombie and Batista-Navarro 2020; Liu et al. 2021; Goffredo et al. 2023a).
@@
  • We train and release RooseBERT, and its training dataset composed of 8K English political debates (5GB).
  • We extensively evaluate the model across four downstream tasks in political discourse analysis: Named Entity Recognition, Sentiment Analysis, Argument Component Detection and Classification, and Argument Relation Prediction and Classification. We compare RooseBERT’s performance with both general-purpose models and other domain-specific PLMs, demonstrating its effectiveness.
@@
By providing a ready-to-use, domain-specific language model, we aim to boost research on this sensible topic, while alleviating the computational costs associated with training such models from scratch.
@@ 2 Related Work
Word embeddings are the foundation of most NLP systems, encoding the semantic and syntactic properties that downstream tasks rely on; in the political domain, well-constructed embeddings are particularly important for capturing domain-specific terminology.
BERT (Devlin et al. 2019) leverages a transformer architecture to produce highly context-sensitive embeddings. It was trained using two different objective learning functions, Masked Language Modelling (MLM) and Next Sentence Prediction (NSP), with the ultimate goal of learning bidirectional contextual representations and capturing sentence-level relationships.
BERT was trained on a large general domain corpora (BooksCorpus and English Wikipedia) and it struggles to capture semantics related to specialised domains. To address this issue, other models have been proposed by pre-training BERT on specific domains, like ClinicalBERT for medical applications (Huang, Altosaar, and Ranganath 2019), SciBERT for scientific texts (Beltagy, Lo, and Cohan 2019), BioBERT for the biomedical domain (Lee et al. 2020), FinBERT for finance (Yang, Uy, and Huang 2020), BERTweet for Twitter data (Nguyen, Vu, and Nguyen 2020) and Mol-BERT (Li and Jiang 2021) for molecular representation.
Different PLMs have been proposed for use in the political domain. Yang et al. (2023) introduced ConfliBERT, a domain-specific PLM for conflict and political violence. ConfliBERT has been trained using 12 datasets and has been evaluated on 18 tasks on which the authors show that the improvement of ConfliBERT is statistically significant in all tasks but three. Kawintiranon and Singh (2022) presented PoliBERTweet, a PLM trained from BERTweet (Nguyen, Vu, and Nguyen 2020) on 83M US 2020 election-related English tweets. Their model outperforms TweetEval (Barbieri et al. 2020) and BERTweet (Nguyen, Vu, and Nguyen 2020) on all tasks related to the political domain than general purpose models.
While ConfliBERT and PoliBERTweet are valuable political-domain models, they are not tailored to the dialogic and argumentative nature of political debates. ConfliBERT focuses on political violence in parliamentary records, while PoliBERTweet targets electoral content on Twitter. Both models lack the turn-taking, rebuttals, and rhetorical complexity central to debates.
@@ 3 Methodology
The American Presidency Project 111www.presidency.ucsb.edu is an authoritative on-line source for presidential public documents. The website includes a list of presidential candidate debates from 1960-2024, including primary and general elections political debates.
@@
The UNGDC (Baturo, Dasandi, and Mikhaylov 2017) is a collection of the United Nations General Assembly debates from 1946 to 2023 with over 10,000 speeches from 202 countries.
House of Commons Parliamentary Debates
@@
The House of Commons Parliamentary Debates corpus (Blumenau 2021) contains data on all parliamentary debates held in the House of Commons between 1979 and 2019. In total, the data includes over 2.5 million speeches drawn from over 50 thousand parliamentary debates.
Australian Parliamentary Debates
@@
The Australia Parliamentary Debates corpus (Katz and Alexander 2023) contains data on the proceedings from each sitting day in the Australian Parliament by the House of Representatives from March 1998 to September 2022.
EUSpeech
@@
The EUSpeech (Schumacher et al. 2016) is a dataset of 18,403 speeches from European Union (EU) leaders (i.e., heads of government in 10 member states, EU commissioners, party leaders in the European Parliament, and European Central Bank and International Monetary Fund leaders) from 2007 to 2015. We only retain English speeches.
ParlEE Plenary Speeches
@@
The ParlEE Plenary Speeches dataset (Sylvester, Greene, and Ebing 2022) is a collection of speeches from eight legislative chambers for Austria, Belgium, France, Germany, Hungary, Ireland, Portugal and the United Kingdom, covering 2009-2019. From this dataset, we only retained the speeches of Ireland and United Kingdom.
ParlScot
@@
The parlScot corpus (Braby and Stewart 2021) contains full-text vectors for all 1.8 million spoken contributions for the Scottish Parliament until 2021, including both speeches in parliament and in committees.
UN Security Council Debates
@@
The UN Security Council Debates Corpus (Schoenfeld et al. 2019) is a dataset of UN Security Council debates between January 1992 and December 2023. In total, the corpus contains 106,302 speeches extracted from 6,233 meeting protocols.
Pre-training Approaches
@@
To effectively pre-train a language model, different approaches can be used (Chalkidis et al. 2020; Yang et al. 2023). One approach is to continue the pre-training (cont) of the original model (BERT) by initialising the model with the original weights and vocabulary and training for additional steps on the domain-specific corpus. The second approach (scr) consists in training BERT from scratch, starting with a random initialisation of its weights and creating a custom vocabulary using the domain-specific corpus.
@@
RooseBERT was trained with both approaches (cont and scr), using cased and uncased vocabularies. When training the scr version of RooseBERT, a custom WordPiece Tokenizer trained on the same training corpus of English political debates was used. Different studies show that scr models gain substantial improvement in performance over cont models (Beltagy, Lo, and Cohan 2019; Gu et al. 2022). This performance gain is largely attributed to the use of a tokenizer with a domain-specific vocabulary, which can encode entire specialised terms as single tokens. In contrast, tokenizers based on general-purpose vocabularies tend to break these terms into multiple sub-tokens, potentially weakening the model’s understanding of domain-specific content. For example, RooseBERT-scr-uncased’s vocabulary shares only 56.32% of its tokens with bert-base-uncased’s vocabulary. Words such as deterrent, endorse, and bureaucrat are represented as single tokens in RooseBERT’s custom vocabulary, while they are absent from BERT’s vocabulary and thus would be split into sub-tokens (see Table 1).
@@
In addition to the cont and scr approaches, each in both cased and uncased variants, RooseBERT was also trained using both base (i.e., BERT-base) and large (i.e., BERT-large) configurations, as larger models have been shown to capture more complex linguistic patterns.
Evaluation Tasks
To evaluate the performance of RooseBERT, we selected four downstream tasks: Named Entity Recognition (NER), Sentiment Analysis, Argument Component Detection and Classification, and Argument Relation Prediction and Classification. These are key tasks in NLP for analysing political debates, and publicly available annotated datasets make it possible to evaluate models on each of them. Accuracy, precision, recall, F1 score and Macro F1 score were used to assess the models.
Named Entity Recognition
The goal of the Named Entity Recognition (NER) task is to identify occurrences of rigid designators in text that belong to predefined semantic categories, such as persons, locations, organisations (Li et al. 2022). We evaluated our model on the political section of the CrossNER dataset (Liu et al. 2021) that contains entity types such as politician, person, organisation, political party, election, country. It is important to note that this section represents only one of the five diverse areas (i.e., politics, natural science, music, literature, artificial intelligence) covered by CrossNER. Although CrossNER-politics is not specifically designed for political debates, we selected it to test the model’s generalizability across different types of political texts, given the scarcity of NER datasets tailored to the politics context.
@@
Sentiment analysis is the task which aims to determine whether the polarity of a text is positive or negative. ParlVote (Abercrombie and Batista-Navarro 2020) is a large annotated corpus of parliamentary debates annotated for the sentiment analysis task. This corpus consists of 34,010 UK parliamentary debates (17,721 annotated as positive, 15,740 as negative) labelled at the speech level.
@@
In Argument Mining (AM), the detection and classification of argument components are challenging tasks that involve extracting argumentative units from natural language text by detecting their boundaries and classifying them as either premises or claims (Lawrence and Reed 2019). In the literature, the most comprehensive resource for political debates annotated with both claims and premises is the ElecDeb60to20 dataset (Goffredo et al. 2023b), a collection of 44 televised U.S. presidential debates from 1960 to 2020. The dataset comprises 38,667 argument components (25,078 claims and 13,589 premises).
@@
In AM, argumentative components, premises and claims, can either support or attack other argument components, forming complex relational structures that are often represented as argumentation graphs (Lawrence and Reed 2019). Identifying these relations is essential for understanding the argumentative structure of a debate. ElecDeb60to20 provides annotations for argument relations. More precisely, 21,689 relations are annotated as support and 3,835 as attack.
@@ 4 Experimental Setup
RooseBERT was trained using both cont and scr approaches, testing both cased and uncased versions with base and large architectures. The base architecture is the same as BERT-Base with 12 layers, 768 hidden units and 12 attention heads. The large model follows BERT-Large, with 24 layers, 1024 hidden units, and 16 attention heads.
@@
Initially, the models were trained to optimise both the Masked Language Modeling (MLM), with a masking probability of 15%, and the Next Sentence Prediction (NSP) objectives. However, as noted in several studies (Liu et al. 2019; Joshi et al. 2020; Yang et al. 2023), incorporating NSP alongside MLM did not lead to significant performance gains compared to training with MLM alone and it only slowed down the training. Therefore, we chose to discard the NSP objective in our training setup. Each pre-training dataset was split into 90% training and 10% testing, with 1% of the test set reserved for perplexity evaluation. These splits were merged to form the final dataset, ensuring balanced representation across sources.
@@
We experimented with a range of hyper-parameters to identify the configuration that most effectively minimised the perplexity and the loss on the testing dataset. Specifically, we tested learning rates of $\{1,3,5\}\times e^{-4}$ and $\{1,3,5\}\times e^{-5}$. The learning rate was linearly increased (warmed up) over the first 10,000 steps, followed by a linear decay for the remainder of training. Once the learning rate was fixed, we experimented with varying number of training steps $\in$ {50K , 100K , 150K , 200K , 250K , 300K}. Following Devlin et al. (2019), each model was trained for 80% of the steps with a maximum sequence length of 128 and with a maximum sequence length of 512 for the remaining 20%. For each training we used eight A100 GPUs and the Transformer Library of HuggingFace (Wolf et al. 2019) with Python 3.10. Each training run was optimised using DeepSpeed (Rasley et al. 2020) ZeRO-2 methods, gradient accumulation, and FP16 precision to reduce training time.
Fine-tuning Setup

### v2

@@ Abstract
The increasing amount of political debates and politics-related discussions calls for the definition of novel computational methods to automatically analyse such content with the final goal of lightening up political deliberation to citizens. However, the specificity of the political language and the argumentative form of these debates (employing hidden communication strategies and leveraging implicit arguments) make this task very challenging, even for current general-purpose pre-trained Language Models. To address this issue, we introduce a novel pre-trained Language Model for political discourse language called RooseBERT. Pre-training a language model on a specialised domain presents different technical and linguistic challenges, requiring extensive computational resources and large-scale data. RooseBERT has been trained on large political debate and speech corpora (8K debates, each composed of several sub-debates on different topics) in English. To evaluate its performances, we fine-tuned it on four downstream tasks related to political debate analysis, i.e., stance detection, sentiment analysis, argument component detection and classification, and argument relation prediction and classification. Our results demonstrate significant improvements over general-purpose Language Models on these four tasks, highlighting how domain-specific pre-training enhances performance in political debate analysis. We release RooseBERT for the research community. [^1] https://github.com/deborahdore/RooseBERT
RooseBERT: A New Deal For Political Language Modelling
Deborah Dore  and Elena Cabrio  and Serena Villata Université Côte d’Azur, CNRS, INRIA, I3S, France {firstname.surname}@univ-cotedazur.fr
@@ 1 Introduction
Political debates, politics-related (online) discussions and opinions represent key elements of public discourse and deliberation, and they are fundamental for citizens to take informed decisions. The ability to automatically analyse such a kind of content is therefore central in Artificial Intelligence, where different aspects of political discourse have been investigated, e.g., performance evaluation (Liu et al., 2024), argumentative structures (Haddadan et al., 2019), sentiment analysis (Abercrombie and Batista-Navarro, 2020), and logical fallacies (Goffredo et al., 2022). These approaches rely on pre-trained Language Models (PLMs) originally developed for general-purpose corpora, e.g., BERT (Devlin et al., 2019). Although these models achieve good results in some of these tasks, they underperform when the goal is to capture the unique features of political content (e.g., domain-specific terminology, implicit argumentation, strategic communication patterns). In the literature, researchers have shown that domain-specific variations in vocabulary and semantics can significantly affect model performance (Nooralahzadeh et al., 2018), avoiding vocabulary mismatches and encoding semantic shifts (Lee et al., 2020).
To address this issue, in this paper, we introduce RooseBERT, a domain-specific language model pre-trained on English political debates. Built on the BERT architecture, widely adopted for domain adaptation tasks (Beltagy et al., 2019; Huang et al., 2019; Lee et al., 2020; Yang et al., 2020; Kawintiranon and Singh, 2022; Yang et al., 2023; Wada et al., 2024; Zeinali et al., 2024), RooseBERT is tailored to capture the distinctive features of political debates. Our aim is to improve the performance of state-of-the-art models on tasks such as sentiment analysis, stance detection, and argument mining which are especially relevant to political analysis. Encoder-based architectures like BERT are especially well-suited for such tasks due to their strength in encoding semantic relationships and contextual information and have proven strong performance on a range of downstream tasks relevant to political discourse analysis (Abercrombie and Batista-Navarro, 2020; Liu et al., 2021; Goffredo et al., 2023a).
@@
  • We train and release RooseBERT, and its training dataset composed of 8K English political debates ($\sim$ 5GB).
  • We extensively evaluate the model across four downstream tasks in political discourse analysis: Stance Detection, Sentiment Analysis, Argument Component Detection and Classification, and Argument Relation Prediction and Classification. We compare RooseBERT’s performance with both general-purpose models and other domain-specific PLMs, demonstrating its effectiveness.
@@
By providing a ready-to-use, domain-specific language model, we aim to boost research on this sensible topic, while alleviating the computational costs of training such models from scratch.
@@ 2 Related Work
Word embeddings are the foundation of most NLP systems, encoding the semantic and syntactic properties that downstream tasks rely on. BERT (Devlin et al., 2019) leverages a transformer architecture to produce highly context-sensitive embeddings. It was trained using two different objective learning functions, Masked Language Modelling (MLM) and Next Sentence Prediction (NSP), with the ultimate goal of learning bidirectional contextual representations and capturing sentence-level relationships. BERT was trained on a large general domain corpora (BooksCorpus and English Wikipedia) and it struggles to capture semantics related to specialised domains. To address this issue, other models have been proposed by pre-training BERT on specific domains, like ClinicalBERT for medical applications (Huang et al., 2019), SciBERT for scientific texts (Beltagy et al., 2019), BioBERT for the biomedical domain (Lee et al., 2020), FinBERT for finance (Yang et al., 2020), BERTweet for Twitter data (Nguyen et al., 2020) and Mol-BERT (Li and Jiang, 2021) for molecular representation.
In the political domain, well-constructed embeddings are particularly important for capturing domain-specific terminology, and different PLMs have been proposed for use in this context. Yang et al. (2023) introduced ConfliBERT, a domain-specific PLM for conflict and political violence. ConfliBERT has been trained using 12 datasets and has been evaluated on 18 tasks on which the authors show that the improvement of ConfliBERT is statistically significant in all tasks but three. Kawintiranon and Singh (2022) presented PoliBERTweet, a PLM trained from BERTweet (Nguyen et al., 2020) on 83M US 2020 election-related English tweets. Their model outperforms TweetEval (Barbieri et al., 2020) and BERTweet (Nguyen et al., 2020) on all tasks related to the political domain than general purpose models. While ConfliBERT and PoliBERTweet are valuable political-domain models, they are not tailored to the dialogic and argumentative nature of political debates. ConfliBERT focuses on political violence in parliamentary records, while PoliBERTweet targets electoral content on Twitter. Both models lack the turn-taking, rebuttals, and rhetorical complexity central to debates.
@@ 3.1 Corpora
The American Presidency Project222www.presidency.ucsb.edu is an authoritative on-line source for presidential public documents. The website includes a list of presidential candidate debates from 1960-2024, including primary and general elections political debates.
@@
is a collection of the UN General Assembly debates (1946- 2023) with over 10k speeches from 202 countries  (Baturo et al., 2017).
The House of Commons Parliamentary Debates
@@
contains data on all parliamentary debates held in the House of Commons between 1979 and 2019. In total, the data includes over 2.5 million speeches drawn from over 50 thousand parliamentary debates (Blumenau, 2021).
The Australian Parliamentary Debates
@@
corpus (Katz and Alexander, 2023) contains data on the proceedings from each sitting day in the Australian Parliament by the House of Representatives from March 1998 to September 2022.
The EUSpeech
@@
(Schumacher et al., 2016) is a dataset of 18,403 speeches from EU leaders (i.e., heads of government in 10 member states, EU commissioners, party leaders in the European Parliament, and European Central Bank and International Monetary Fund leaders) from 2007 to 2015.
The ParlEE Plenary Speeches
@@
dataset (Sylvester et al., 2022) is a collection of speeches from eight legislative European chambers (Austria, Belgium, France, Germany, Hungary, Ireland, Portugal and the United Kingdom) covering 2009-2019. Only Irish legislative chamber speeches were kept.
The ParlScot
@@
corpus (Braby and Stewart, 2021) contains full-text vectors for 1.8 million spoken contributions for the Scottish Parliament until 2021 (speeches in both parliament and in committees).
The UN Security Council Debates
@@
corpus (Schoenfeld et al., 2019) is a dataset of UN Security Council debates between January 1992 and December 2023. The corpus contains 106,302 speeches extracted from 6,233 meeting protocols.
3.2 Pre-training Approaches
@@
To effectively pre-train a language model, different approaches can be used (Chalkidis et al., 2020; Yang et al., 2023). One approach is to continue the pre-training (cont) of the original model (BERT) by initialising the model with the original weights and vocabulary and training for additional steps on the domain-specific corpus. The second approach (scr) consists in training BERT from scratch, starting with a random initialisation of its weights and creating a custom vocabulary using the domain-specific corpus.
@@
RooseBERT was trained with both approaches (cont and scr), using cased and uncased vocabularies. When training the scr version of RooseBERT, a custom WordPiece Tokenizer trained on the same training corpus of English political debates was used. Different studies show that scr models gain substantial improvement in performance over cont models (Beltagy et al., 2019; Gu et al., 2022). This performance gain is largely attributed to the use of a tokenizer with a domain-specific vocabulary, which can encode entire specialised terms as single tokens. In contrast, tokenizers based on general-purpose vocabularies tend to break these terms into multiple sub-tokens, potentially weakening the model’s understanding of domain-specific content. For example, RooseBERT-scr-uncased’s vocabulary shares only 56.32% of its tokens with bert-base-uncased’s vocabulary. Words such as deterrent, endorse, and bureaucrat are represented as single tokens in RooseBERT’s custom vocabulary, while they are absent from BERT’s vocabulary and thus would be split into sub-tokens (see Table 1).
@@
3.3 Evaluation Tasks
To evaluate the performance of RooseBERT, we selected four downstream tasks: Stance Detection, Sentiment Analysis, Argument Component Detection and Classification, and Argument Relation Prediction and Classification. These tasks are central to NLP research on political debates, and publicly available annotated datasets enable systematic model evaluation. However, due to the limited availability of datasets focused specifically on political debates, we also included a dataset (for stance detection) that is not political-oriented. Accuracy was used as the evaluation metric for Sentiment Analysis, while Average Macro F1 was employed for the remaining tasks, in line with the metrics adopted by the original dataset authors.
@@ 3.3 Evaluation Tasks
Sentiment analysis is the task which aims to determine whether the polarity of a text is positive or negative. ParlVote (Abercrombie and Batista-Navarro, 2020) is a large annotated corpus of parliamentary debates annotated for the sentiment analysis task. This corpus consists of 34,010 UK parliamentary debates (17,721 annotated as positive, 15,740 as negative) labelled at the speech level.
Stance Detection
Stance detection is the task of automatically identifying positions on a specific topic in text (Mohammad et al., 2017). The VAST dataset (Allaway and McKeown, 2020) is an annotated dataset collected from The New York Times Room for Debate section, part of the Argument Reasoning Comprehension (ARC) Corpus (Habernal et al., 2017). The VAST dataset comprises 23,525 sentences across 5,634 topics, each annotated as pro (supporting the topic), con (opposing the topic), or neutral.
@@
In Argument Mining (AM), the detection and classification of argument components are challenging tasks that involve extracting argumentative units from natural language text by detecting their boundaries and classifying them as either premises or claims (Lawrence and Reed, 2019). In the literature, the most comprehensive resource for political debates annotated with both claims and premises is the ElecDeb60to20 dataset (Goffredo et al., 2023b), a collection of 44 televised U.S. presidential debates from 1960 to 2020. The dataset comprises 38,667 argument components (25,078 claims and 13,589 premises).
@@
In AM, argumentative components, i.e., premises and claims, can either support or attack other argument components, forming complex relational structures that are often represented as argumentation graphs (Lawrence and Reed, 2019). Identifying these relations is essential for understanding the argumentative structure of a debate. ElecDeb60to20 provides annotations for argument relations. More precisely, 21,689 relations are annotated as support and 3,835 as attack.
@@ 4.1 Pre-training Setup
RooseBERT was trained using both cont and scr approaches, in both cased and uncased versions. The architecture is the same as BERT-Base with 12 layers, 768 hidden units and 12 attention heads. See Appendix A for the large model training.
@@
Initially, the models were trained to optimise both the Masked Language Modeling (MLM), with a masking probability of 15%, and the Next Sentence Prediction (NSP) objectives. However, as noted in several studies (Liu et al., 2019; Joshi et al., 2020; Yang et al., 2023), incorporating NSP alongside MLM did not lead to significant performance gains compared to training with MLM alone and it only slowed down the training. Therefore, we chose to discard the NSP objective in our training setup. Each pre-training dataset was split into 95% training and 5% testing, with the test set also reserved for perplexity evaluation. These splits were merged to form the final dataset, ensuring balanced representation across sources.
@@
We experimented with a range of hyper-parameters to identify the configuration that most effectively minimised the perplexity and the loss on the testing dataset. Specifically, we tested learning rates of $\{1,3,5\}\times e^{-4}$ and $\{1,3,5\}\times e^{-5}$. The learning rate was linearly increased (warmed up) over the first 10,000 steps, followed by a linear decay for the remainder of training. Once the learning rate was fixed, we experimented with varying number of training steps $\in$ {50K , 100K , 150K , 200K , 250K , 300K}. Following Devlin et al. (2019), each model was trained for 80% of the steps with a maximum sequence length of 128 and with a maximum sequence length of 512 for the remaining 20%. For each training we used eight A100 GPUs and the Transformer Library of HuggingFace (Wolf et al., 2019) with Python 3.10. Each training run was optimised using DeepSpeed (Rasley et al., 2020) ZeRO-2 methods, gradient accumulation, and FP16 precision to reduce training time.
4.2 Fine-tuning Setup